OpenStudy (anonymous):

Write an equation for a line perpendicular to and passing through the point (-6,2)

OpenStudy (anonymous):

perpendicular to what?

OpenStudy (anonymous):

perpendicular to y=-2x-2

OpenStudy (anonymous):

perpendicular to \(y=-2x-2\) a line with slope \(-2\) will have slope \(\frac{1}{2}\)

OpenStudy (anonymous):

your slope is \(\frac{1}{2}\) and your point is \((-6,2) \) use the point slope formula to find the equation of the line

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] with \(m=\frac{1}{2},x_1=-6,y_1=2\) \[y-2=\frac{1}{2}(x+6)\]

OpenStudy (anonymous):

If an equation is in y = mx+b form, then the slope is m. The slope of a perpendicular has the opposite reciprocal as it's slope. That is, flip it and then make it negative.

OpenStudy (anonymous):

\[y-2=\frac{1}{2}x+3\] \[y=\frac{1}{2}x+5\]

OpenStudy (anonymous):

Thank you! I have another problem if you're willing to help!(:

OpenStudy (anonymous):

It costs XYZ Corporation $7500 to produce 1400 units and $4440 to produce 500 units. Let C represent the cost in dollars and x represent the number of units produced. Find a linear equation for C in terms of x.